makes on angle of θ5.00 oits direction of notion Determne its Speed (a) at this manant, and (b) 2cos akr aueeraton 1.05m/s2 (0335:

please answer !
Answer #1

a = acceleration = 1.05 m/s2

\theta = angle of acceleration with the direction of motion or the tangent to circle = 25 deg

ar = radial acceleration acting towards the center

at = tangential acceleration acting parallel to tangent

Radial acceleration is given as

ar = a Sin\theta

ar = 1.05 Sin25 = 0.44 m/s2

tangential acceleration is given as

at = a Cos\theta

at = 1.05 Cos25 = 0.95 m/s2

v = speed

radial acceleration is given as

ar = v2/r

0.44 = v2/1.95

v = 0.93 m/s

b)

vo = intial velocity = 0.93 m/s

v = final velocity after time "t"

a = acceleration = 0.95 m/s2

t = time

using the equation

v = vo + a t

v = 0.93 + (0.95) (2)

v = 2.83 m/s